# Can Prompt Difficulty be Online Predicted for Accelerating RL Finetuning of Reasoning Models?

### v4

@@ 3.2 Bayesian Inference towards Prompt Success Rate
$\gamma^{0}_{\tau}\sim\mathrm{Beta}(\alpha_{\tau}^{0},\beta_{\tau}^{0}),$ (10)
@@
$\underbrace{p(\gamma^{t}_{\tau}\mid H_{t})}_{\text{{Updated Posterior}}}\propto\underbrace{p(\bm{r}^{t}_{\tau}\mid\gamma^{t}_{\tau})}_{\text{{Likelihood}}}\cdot\underbrace{p(\gamma^{t}_{\tau}\mid H_{t-1})}_{\text{{Conjugate Prior}}},$ (11)
@@
$\gamma^{t}_{\tau}\mid H_{t}\sim\mathrm{Beta}(\alpha^{t^{\prime}}_{\tau},\beta^{t^{\prime}}_{\tau}),$ (12)
@@
$\alpha^{t^{\prime}}_{\tau}=\alpha^{t}_{\tau}+s^{t}_{\tau},\quad\beta^{t^{\prime}}_{\tau}=\beta^{t}_{\tau}+k-s^{t}_{\tau}.$ (13)
@@
$\alpha^{t+1}_{\tau}=\alpha^{t^{\prime}}_{\tau},\quad\beta^{t+1}_{\tau}=\beta^{t^{\prime}}_{\tau}.$ (14)
@@
$\alpha^{t^{\prime}}_{\tau}=\lambda\cdot\alpha^{t}_{\tau}+(1-\lambda)\cdot\alpha_{\tau}^{0}+s^{t}_{\tau},\quad\beta^{t^{\prime}}_{\tau}=\lambda\cdot\beta^{t}_{\tau}+(1-\lambda)\cdot\beta_{\tau}^{0}+k-s^{t}_{\tau}.$ (15)
@@
Figure 4: Framework Overview. Left: Comparison between Dynamic Sampling (Oracle), which filters prompts based on actual LLM evaluation on candidates, and our Model Predictive Prompt Selection (MoPPS), which predicts success rates to avoid extra inference cost. Right: MoPPS predicts success rates for candidates from posterior parameters, based on which prompts closest to a target $\gamma^{*}$ are selected for training; the posterior is then updated using new feedback.
3.3 Model Predictive Prompt Selection
@@
$\hat{\gamma}^{t}_{\tau}\sim\mathrm{Beta}(\alpha_{\tau}^{t},\beta_{\tau}^{t})\quad\forall\tau\in\mathcal{T}_{t}^{\hat{\mathcal{B}}}.$ (16)
@@
Importantly, we adopt Thompson Sampling in this work for its simplicity and natural exploration-exploitation trade-off, though MoPPS can also be seamlessly combined with other acquisition strategies such as UCB (Auer, 2002). Besides, our lightweight prediction allows extending $\mathcal{T}_{t}^{\hat{\mathcal{B}}}$ to the entire pool $\mathcal{T}$ at negligible cost, which is infeasible for prior methods that rely on exact evaluation.
@@
Prior works (Bae et al., 2025; Chen et al., 2025) indicate that prompts with intermediate difficulty, i.e., success rates near a target value $\gamma^{*}$, typically around $0.5$, yield the most informative gradients for RL finetuning. Leveraging this insight, at each step, we construct the training batch $\mathcal{T}_{t}^{\mathcal{B}}$ by selecting the $\mathcal{B}$ prompts whose sampled success rates $\hat{\gamma}^{t}_{\tau}$ are closest to $\gamma^{*}$:
@@ 3.4 Implementation Pipeline
$\text{Model Predictive Prompt Selection:}\quad\mathcal{T}_{t}^{\hat{\mathcal{B}}}\xrightarrow{\text{{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}Predict}}}}\{\mathcal{T}_{t}^{\hat{\mathcal{B}}},\hat{\Gamma}_{t}^{\hat{\mathcal{B}}}\}\xrightarrow{\text{{Select}}}\{\mathcal{T}_{t}^{\mathcal{B}}\}.$ (18)
@@
   Randomly sample a candidate set $\mathcal{T}_{t}^{\hat{\mathcal{B}}}=\{\hat{\tau}_{t,i}\}_{i=1}^{\hat{\mathcal{B}}}$ from $\mathcal{T}$;
@@
4 Experiments
This section includes comprehensive experiments to evaluate that MoPPS enables accurate online prediction of prompt difficulty, accelerates RL finetuning, and improves performance compared to baselines. Further analyses confirm its flexibility across selection strategies, benefit from prior knowledge, necessity of online posterior updates, and compatibility with various RL algorithms.
4.1 Experimental Setup
@@
Figure 5: Training curves of MoPPS and baseline methods across three reasoning tasks with varying backbone sizes. Notably, DS serves as an oracle baseline, as it relies on expensive exact LLM evaluations and demands significantly more rollouts.
Reasoning Tasks:
  1. 1.
    Mathematics: We train LLMs on the training split of the MATH dataset (Hendrycks et al., 2021), which consists of problems from mathematics competitions. Following prior work (Luo et al., 2025b), we use the DeepSeek-R1 distillation models DeepSeek-R1-Distill-Qwen-1.5B and DeepSeek-R1-Distill-Qwen-7B (Guo et al., 2025), and track performance on AIME24 during training. Final evaluations are conducted on benchmarks including AIME24, AMC23, MATH500 (Lightman et al., 2023), Minerva Math (Minerva.) (Lewkowycz et al., 2022), and OlympiadBench (Olympiad.) (He et al., 2024).
  2. 2.
    Planning: We adopt the Countdown Number Game, which requires combining given numbers using basic arithmetic operations to reach a target value. Training is performed on a subset of the Countdown-34 dataset (Pan et al., 2025), with performance tracked on a held-out split. Final evaluation is conducted on both Countdown-34 (CD-34) and a more challenging variant, Countdown-4 (CD-4). Following Chen et al. (2025), we use two base models: Qwen2.5-3B and Qwen2.5-7B (Yang et al., 2024a).
  3. 3.
    Visual Geometry: Geometry problems require both visual understanding and reasoning. Two vision-language models, Qwen2.5-VL-3B-Instruct and Qwen2.5-VL-7B-Instruct (Bai et al., 2025), are trained on the training split of the Geometry3k dataset (Lu et al., 2021; Hiyouga, 2025) and evaluated on its test split.
@@
Two common sampling strategies are compared with MoPPS: (1) Uniform, which samples prompts uniformly from the prompt pool; and (2) Dynamic Sampling (DS), which oversamples prompts and filters out uninformative ones based on their exact evaluation, as described in Eq. (6). Notably, DS serves as an oracle baseline since it relies on real evaluation feedback. Our focus is on reducing computational overhead relative to DS rather than outperforming it in overall accuracy.
4.2 Main Results
4.2.1 Highly Correlated Difficulty Prediction
@@
$\rho=\frac{\mathrm{cov}(\mathrm{rank}(\hat{\Gamma}^{\mathcal{B}}),\mathrm{rank}(\widetilde{\Gamma}^{\mathcal{B}}))}{\sigma_{\mathrm{rank}(\hat{\Gamma}^{\mathcal{B}})}\cdot\sigma_{\mathrm{rank}(\widetilde{\Gamma}^{\mathcal{B}})}}$, (19)
@@
In Fig. 1, our method exhibits consistently high rank correlation ($\rho>0.5$) between the estimated difficulty and the ground-truth with extremely low $p$-values across reasoning tasks and diverse backbone models. Besides, a clear training progresses can be observed that the correlation steadily improves until stabilizing at a high level. This validates that our Bayesian surrogate accumulates meaningful evidence over time, progressively refining its belief about prompt difficulty.
@@
Table 1: Evaluation across mathematics benchmarks. ‘+’ indicates finetuning with the corresponding method. Accuracy is computed as the average pass@1 over 16 independent generations per problem. ‘Avg.’ denotes average accuracy across benchmarks, and ‘Rollouts’ indicates the number of rollout samples during finetuning. Bold indicates the best result; underlined indicates the second best. ‘MoPPS w/ prior’ means incorporating prior knowledge into MoPPS.
@@
4.2.2 Accelerated RL Finetuning
We compare the training performance of MoPPS with baselines across different scenarios and backbone models. Fig. 5 shows the training curves, and Table 1 summarizes the final evaluation results on the mathematics task. Thanks to reliable difficulty prediction, the proposed MoPPS method achieves both training acceleration and better final performance compared to uniform prompt selection. On MATH, uniform selection suffers from performance collapse, likely due to entropy collapse (Liu et al., 2025a). Effective online prompt selection methods can mitigate this issue and sustain continuous progress, with MoPPS achieving approximately $\frac{32.92-26.46}{26.46}\approx\bm{24.4\%}$ relative improvement on AIME24 and about $\bm{6.7\%}$ average relative improvement on multiple benchmarks with the 1.5B backbone compared to Uniform. On Countdown and Geometry, MoPPS consistently accelerates training across backbone sizes, reaching nearly $\bm{1.8\times}$ speedup. Compared to DS, MoPPS attains comparable performance with only $\bm{25\%}$ rollouts on MATH and $\bm{21\%}$ on Countdown (Table 3). This efficiency gain stems from DS’s requirement to evaluate a larger candidate prompt set using costly LLM inference at each step, whereas MoPPS amortizes this via posterior sampling.
@@ 4.2.2 Accelerated RL Finetuning
Figure 6: Ablation studies on selection strategies, prior knowledge, and online posterior updates conducted on the MATH task using DeepSeek-R1-Distill-Qwen-1.5B. (a) Comparison of $\text{Top-}\mathcal{B}$ and Threshold selection strategies, with and without prior knowledge. (b) Training performance of offline selection (prior only) versus MoPPS with prior and online posterior updates. (c) Spearman rank correlation over training steps for offline variant and MoPPS with prior.
4.3 Additional Analysis
4.3.1 Algorithm Compatibility
Table 2: Evaluation on Countdown with PPO and Reinforce++. MoPPS consistently improves pass@1 accuracy on CD-34 and harder CD-4.
@@ 4.3.1 Algorithm Compatibility
4.3.2 The Effects of Prior Knowledge and Selection Strategies
Our default implementation adopts a uniform prior, $\text{Beta}(1,1)$, and a $\text{Top-}\mathcal{B}$ selection strategy. To examine the flexibility of our method, we evaluate (i) an alternative ‘Threshold’ selection strategy, as used in Bae et al. (2025), which samples prompts with predicted success rates falling within a fixed interval, i.e., $\gamma_{min}\leq\hat{\gamma}_{\tau}\leq\gamma_{max}$, and (ii) the integration of prior knowledge by pre-evaluating all prompts using the base model and initializing the Beta parameters $\{\alpha,\beta\}$ accordingly. As shown in Fig. 6(a), both strategies improve over uniform selection, and $\text{Top-}\mathcal{B}$ performs better. Incorporating prior knowledge further enhances training efficiency, though our method remains effective even without such prior.
4.3.3 Role of Online Posterior Updates
To assess the importance of online posterior updates, we consider an ‘Offline’ variant that uses only prior knowledge for prompt selection without updating the posterior during training. As shown in Fig. 6(b), while the offline variant benefits from a strong initialization and achieves reasonable early-stage performance, its accuracy degrades in later stages. This degradation arises from the its inability to adapt to the evolving policy, resulting in outdated difficulty estimates and increasingly suboptimal prompt selection. As evidenced in Fig. 6(c), the offline variant suffers a decline in correlation over time, whereas the online variant improves continually by updating its posterior with new feedback.
@@ 5 Related Works
Reinforcement learning has emerged as a powerful paradigm for aligning LLMs with desired behaviors. Reinforcement Learning with Human Feedback (RLHF), has demonstrated remarkable success in improving instruction-following capabilities and ensuring the safety of LLMs (Dong et al., 2024; Dai et al., 2023; Sun et al., 2023; Zheng et al., 2023). More recently, advances show that Reinforcement Learning with Verifiable Rewards (RLVR) (Jaech et al., 2024; Guo et al., 2025; Team et al., 2025; Chu et al., 2025; Pan et al., 2025) significantly enhances LLMs’ reasoning capabilities in structured domains, such as mathematics, where reward signals can be automatically verified. Among RL algorithms, PPO (Schulman et al., 2017) remains a widely adopted method. GRPO (Shao et al., 2024) has gained traction more recently as it eliminates the computationally expensive value network in PPO by estimating advantages using a lightweight group-normalized manner. Several recent works further improve these algorithms by avoiding bias and training collapse, reducing overhead, and enhancing sample efficiency (Yuan et al., 2025; Yue et al., 2025; Liu et al., 2025b; Yu et al., 2025; Kazemnejad et al., 2024; Hu, 2025). Moreover, many efforts push the performance frontier across diverse domains and model scales (Luo et al., 2025b; Dang & Ngo, 2025; Luo et al., 2025a; Zeng et al., 2025; Meng et al., 2025; Xu et al., 2024), while others provide infrastructure for scalable RL-based LLM training (Sheng et al., 2024).

### v5

@@ 3.2. Bayesian Inference towards Prompt Success Rate
Figure 3. Framework Overview. Left: Comparison between Dynamic Sampling (Oracle), which filters prompts based on actual LLM evaluation on candidates, and our Model Predictive Prompt Selection (MoPPS), which predicts success rates to avoid extra inference cost. Right: MoPPS predicts success rates for candidates from posterior parameters, based on which prompts closest to a target $\gamma^{*}$ are selected for training; the posterior is then updated using new feedback.
[…]
@@
(10) $\gamma^{0}_{\tau}\sim\mathrm{Beta}(\alpha_{\tau}^{0},\beta_{\tau}^{0}),$
@@
(11) $\underbrace{p(\gamma^{t}_{\tau}\mid H_{t})}_{\text{{Updated Posterior}}}\propto\quad\underbrace{p(\bm{r}^{t}_{\tau}\mid\gamma^{t}_{\tau})}_{\text{{Likelihood}}}\cdot\underbrace{p(\gamma^{t}_{\tau}\mid H_{t-1})}_{\text{{Conjugate Prior}}},$
@@
(12) $\gamma^{t}_{\tau}\mid H_{t}\sim\mathrm{Beta}(\alpha^{t^{\prime}}_{\tau},\beta^{t^{\prime}}_{\tau}),$
@@
(13) $\alpha^{t^{\prime}}_{\tau}=\alpha^{t}_{\tau}+s^{t}_{\tau},\quad\beta^{t^{\prime}}_{\tau}=\beta^{t}_{\tau}+k-s^{t}_{\tau}.$
@@
(14) $\alpha^{t+1}_{\tau}=\alpha^{t^{\prime}}_{\tau},\quad\beta^{t+1}_{\tau}=\beta^{t^{\prime}}_{\tau}.$
@@
(15) $\alpha^{t^{\prime}}_{\tau}=\lambda\cdot\alpha^{t}_{\tau}+(1-\lambda)\cdot\alpha_{\tau}^{0}+s^{t}_{\tau},$
$\beta^{t^{\prime}}_{\tau}=\lambda\cdot\beta^{t}_{\tau}+(1-\lambda)\cdot\beta_{\tau}^{0}+k-s^{t}_{\tau}.$
@@
3.3. Model Predictive Prompt Selection
@@
(16) $\hat{\gamma}^{t}_{\tau}\sim\mathrm{Beta}(\alpha_{\tau}^{t},\beta_{\tau}^{t})\quad\forall\tau\in\mathcal{T}_{t}^{\hat{\mathcal{B}}}.$
@@
Importantly, we adopt Thompson Sampling in this work for its simplicity and natural exploration-exploitation trade-off, which inherently serves as an uncertainty-aware data curation mechanism. MoPPS can also be seamlessly combined with other acquisition strategies such as UCB (Auer, 2002). In addition, our lightweight prediction allows us to extend $\mathcal{T}_{t}^{\hat{\mathcal{B}}}$ to the entire pool $\mathcal{T}$ at negligible cost, which significantly improves the exploration space and is infeasible for prior methods that rely on exact evaluation.
@@
Prior works (Bae et al., 2025; Chen et al., 2025) indicate that prompts with intermediate difficulty, i.e., success rates near a target value $\gamma^{*}$, typically around $0.5$, yield the most informative gradients for RL finetuning. Viewed from the perspective of curriculum learning (Bengio et al., 2009), maintaining a target difficulty induces an implicit curriculum that gradually transitions from easier to harder prompts as the model improves. Leveraging this insight, at each step, we construct the training batch $\mathcal{T}_{t}^{\mathcal{B}}$ by selecting the $\mathcal{B}$ prompts whose sampled success rates $\hat{\gamma}^{t}_{\tau}$ are closest to $\gamma^{*}$:
@@ 3.4. Implementation Pipeline
(18) Model Predic tive Prompt Selection:
$\mathcal{T}_{t}^{\hat{\mathcal{B}}}\xrightarrow{\text{{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}Predict}}}}\{\mathcal{T}_{t}^{\hat{\mathcal{B}}},\hat{\Gamma}_{t}^{\hat{\mathcal{B}}}\}\xrightarrow{\text{{Select}}}\{\mathcal{T}_{t}^{\mathcal{B}}\}.$
@@
   Randomly sample candidate set $\mathcal{T}_{t}^{\hat{\mathcal{B}}}=\{\hat{\tau}_{t,i}\}_{i=1}^{\hat{\mathcal{B}}}$ from $\mathcal{T}$;
@@
4. Experiments
This section conducts experiments to evaluate whether MoPPS can online predict prompt difficulty, accelerate RL finetuning, and improve performance. Additional analyses explore its flexibility across selection strategies, use of prior knowledge, need for posterior updates, and compatibility with various RL algorithms.
4.1. Experimental Setup
@@
Mathematics Task
We train LLMs on the training split of the MATH dataset (Hendrycks et al., 2021), which consists of problems from mathematics competitions. Following prior work (Luo et al., 2025b), we use the DeepSeek-R1 distillation models R1-Distill-Qwen-1.5B and R1-Distill-Qwen-7B (Guo et al., 2025), and track performance on AIME24 during training. Final evaluations are conducted on benchmarks including AIME24, AMC23, MATH500 (Lightman et al., 2023), Minerva Math (Minerva.) (Lewkowycz et al., 2022), and OlympiadBench (Olympiad.) (He et al., 2024).
Planning Task
We adopt the Countdown Number Game, which requires combining given numbers using basic arithmetic operations to reach a target value. Training is performed on a subset of the Countdown-34 (CD-34) dataset (Pan et al., 2025), with performance tracked on a held-out split. Final evaluation is conducted on both CD-34 and a more challenging variant, Countdown-4 (CD-4). Following (Chen et al., 2025), we use two base models: Qwen2.5-3B and Qwen2.5-7B (Yang et al., 2024a).
Visual Geometry Task
Geometry problems require both visual understanding and reasoning. Two vision language models, Qwen2.5-VL-3B-Instruct and Qwen2.5-VL-7B-Instruct (Bai et al., 2025), are trained on the training split of the Geometry3k dataset (Lu et al., 2021; Hiyouga, 2025) and evaluated on its test split.
@@
Two common sampling strategies are compared with MoPPS: (1) Uniform, which samples prompts uniformly from the prompt pool; (2) History Resampling (HR) (Zhang et al., 2025), which excludes fully solvable prompts each epoch; and (3) Dynamic Sampling (DS) (Yu et al., 2025), which oversamples prompts and filters out uninformative ones based on their exact evaluation, as described in Eq. (6). Notably, DS serves as an Oracle and latest SOTA baseline since it relies on real evaluation feedback. Our focus is on reducing computational overhead relative to DS rather than outperforming it in overall accuracy.
4.2. Main Results
4.2.1. Highly Correlated Difficulty Prediction
Figure 4. Spearman rank correlation and $p$-value over training steps between the predicted prompt difficulty from our Bayesian surrogate and the empirical success rate. The strong correlation indicates that our method effectively predicts prompt difficulty without incurring costly LLM inferences.
Spearman rank correlation and \$p\$-value over training steps between the predicted prompt difficulty from our Bayesian surrogate and the empirical success rate. The strong correlation indicates that our method effectively predicts prompt difficulty without incurring costly LLM inferences.
@@ 4.2.1. Highly Correlated Difficulty Prediction
(19) $\rho=\frac{\mathrm{cov}(\mathrm{rank}(\hat{\Gamma}^{\mathcal{B}}),\mathrm{rank}(\widetilde{\Gamma}^{\mathcal{B}}))}{\sigma_{\mathrm{rank}(\hat{\Gamma}^{\mathcal{B}})}\cdot\sigma_{\mathrm{rank}(\widetilde{\Gamma}^{\mathcal{B}})}}$,
@@
In Fig. 4, MoPPS exhibits consistently high rank correlation ($\rho>0.5$) between the estimated difficulty and the ground-truth with extremely low $p$-values across reasoning tasks and diverse backbone models. Besides, a clear training progresses can be observed that the correlation steadily improves until stabilizing at a high level. This validates that our Bayesian surrogate accumulates meaningful evidence over time, progressively refining its belief about prompt difficulty.
@@
Figure 5. Training curves of MoPPS and baselines across three reasoning tasks with varying backbone sizes. Notably, DS serves as an oracle baseline, as it relies on expensive exact LLM evaluations and demands significantly more rollouts.
Table 1. Evaluation across mathematics benchmarks. ‘+’ indicates finetuning with the corresponding method. Accuracy is computed as the average pass@1 over 16 independent generations per problem. ‘Avg.’ denotes average accuracy across benchmarks, and ‘Rollouts’ indicates the number of rollout samples during finetuning. Bold indicates the best result; underlined indicates the second best.
@@
4.2.2. Accelerated RL Finetuning
We compare the training performance of MoPPS with baselines across different scenarios and backbone models. Fig. 5 shows the training curves, and Table 1 summarizes the final evaluation results on the mathematics task. Thanks to reliable difficulty prediction, the proposed MoPPS method achieves both training acceleration and better final performance compared to uniform prompt selection. On MATH, uniform selection suffers from performance collapse, likely due to entropy collapse (Liu et al., 2025a). Effective online prompt selection methods can mitigate this issue and sustain continuous progress, with MoPPS achieving approximately $\frac{32.92-26.46}{26.46}\approx\bm{24.4\%}$ relative improvement on AIME24 and about $\bm{6.7\%}$ average relative improvement on multiple benchmarks with the 1.5B backbone compared to Uniform. On Countdown and Geometry, MoPPS consistently accelerates training across backbone sizes, reaching nearly $\bm{1.8\times}$ speedup. Compared to DS, MoPPS attains comparable performance with only $\bm{25\%}$ rollouts on MATH (Table 1) and $\bm{21\%}$ on Countdown (Table 3). This efficiency gain stems from DS’s requirement to evaluate a larger candidate prompt set using costly LLM inference at each step, whereas MoPPS amortizes this via lightweight model prediction.
@@ 4.2.2. Accelerated RL Finetuning
4.3. Additional Analysis
4.3.1. Algorithm Compatibility
4.3.2. The Effects of Prior Knowledge and Selection Strategies
Our default implementation adopts a uniform prior, $\text{Beta}(1,1)$, and a $\text{Top-}\mathcal{B}$ selection strategy. To examine the flexibility of our method, we evaluate (i) an alternative Threshold selection strategy, as used in (Bae et al., 2025), which samples prompts with predicted success rates falling within a fixed interval, i.e., $\gamma_{min}\leq\hat{\gamma}_{\tau}\leq\gamma_{max}$, and (ii) the integration of prior knowledge (‘w/ prior’) by pre-evaluating all prompts using the base model and initializing the Beta parameters $\{\alpha,\beta\}$ accordingly. As shown in Fig. 6(a), both strategies improve over uniform selection, and $\text{Top-}\mathcal{B}$ performs better. Incorporating prior knowledge further enhances training efficiency, though our method remains effective even without such prior.
4.3.3. Role of Online Posterior Updates
To assess the importance of online posterior updates, we consider an Offline variant that uses only prior knowledge for prompt selection without updating the posterior during training. As shown in Fig. 6(b), the offline variant benefits from strong initialization and performs competitively in early stages, but its accuracy degrades later. This degradation arises from the its inability to adapt to the evolving policy, resulting in outdated difficulty estimates and increasingly suboptimal prompt selection. As evidenced in Fig. 6(c), the offline variant suffers a decline in correlation over time, whereas the online variant improves continually by updating its posterior with new feedback.
Table 2. Evaluation on Countdown with PPO and Reinforce++ using Qwen2.5-3B. MoPPS consistently improves pass@1 accuracy on CD-34 and the harder CD-4 benchmark compared to uniform selection.
@@ 4.3.3. Role of Online Posterior Updates
Figure 6. Ablation studies on selection strategies, prior knowledge, and online posterior updates on MATH using R1-Distill-Qwen-1.5B. (a) Comparison of $\text{Top-}\mathcal{B}$ and Threshold selection strategies, with and without prior knowledge. (b) Training performance of offline selection (prior only) versus MoPPS with prior and online posterior updates. (c) Spearman rank correlation over training steps for offline variant and MoPPS with prior.
@@ 5. Related Works
Reinforcement learning has emerged as a powerful paradigm for aligning LLMs with desired behaviors. Reinforcement Learning with Human Feedback (RLHF), has demonstrated remarkable success in improving instruction-following capabilities and ensuring the safety of LLMs (Dong et al., 2024; Dai et al., 2023; Sun et al., 2023; Zheng et al., 2023). More recently, advances show that Reinforcement Learning with Verifiable Rewards (RLVR) (Jaech et al., 2024; Guo et al., 2025; Team et al., 2025; Chu et al., 2025; Pan et al., 2025) significantly enhances LLMs’ reasoning capabilities in structured domains, such as mathematics, where reward signals can be automatically verified. Among RL algorithms, PPO (Schulman et al., 2017) remains a widely adopted method. GRPO (Shao et al., 2024) has gained traction more recently as it eliminates the computationally expensive value network in PPO by estimating advantages using a lightweight group-normalized manner. Several recent works further improve these algorithms by avoiding bias and training collapse, reducing overhead, and enhancing sample efficiency (Yuan et al., 2025; Yue et al., 2025; Liu et al., 2025b; Yu et al., 2025; Kazemnejad et al., 2024; Hu, 2025). Moreover, many efforts push the performance frontier across diverse domains and model scales (Luo et al., 2025b; Dang and Ngo, 2025; Luo et al., 2025a; Zeng et al., 2025; Meng et al., 2025; Xu et al., 2024), while others provide infrastructure for scalable RL-based LLM training (Sheng et al., 2024).